OpenStudy (anonymous):

Simplify log 2 + log 3 + log 4. A. log 9 B. log 24 C. 3 D. 4

terenzreignz (terenzreignz):

use this property \[\huge \log(M) + \log(N) = \log(MN)\]

OpenStudy (anonymous):

Okay I got 3

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

No it is log(4*3*2) which is log(24)

OpenStudy (anonymous):

Express 62 = 36 using logarithmic notation. A. 2 = log6 36 B. 6 = log2 36 C. 36 = log2 6 D. 6 = log36 2

OpenStudy (anonymous):

option A is correct!